OpenStudy (anonymous):

What is the vertex of the graph of the function y= |-3x + 2| Please somebody help I will reward!

OpenStudy (mathmath333):

x-coordinate is what makes what is inside the absolute value =0, then the y-coordinate is the value you get when you put that in \(\large\tt \begin{align} \color{black}{ y=\mid -3x+2\mid\hspace{.33em}\\~\\ 0=\mid -3x+2\mid\hspace{.33em}\\~\\ \fbox {x=2/3} \hspace{.33em}\\~\\~\\ y=\mid -3x+2\mid\hspace{.33em}\\~\\ y=\mid -3\times \dfrac{2}{3}+2\mid\hspace{.33em}\\~\\ \fbox{y=0}\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

https://www.desmos.com/calculator/tto25jv8kd